# Large Language Model Adaptation for Networking

### v1

@@ 2.3. Domain-Adapted LLMs
The impressive performance of LLMs has sparked pioneering research to adapt LLMs for other domains (Driess et al., 2023; Lin et al., 2023; Liu et al., 2023a; Zitkovich et al., 2023). For example, PaLM-E (Driess et al., 2023) adapts PaLM (Chowdhery et al., 2023) to generate step-by-step controlling commands for robotic manipulation. ESMFFold (Lin et al., 2023) showcases the successful applications of LLM in biological fields, which leverages LLM to accurately predict atomic-level protein structure.
Inspired by these promising outcomes, this work explores the adaptation of LLMs for networking to address the limitations of existing learning-based algorithms, hoping to pave the way for more sustainable design philosophy of networking solutions. Unfortunately, although some prior studies have showcased that LLMs are capable to generate some technical documents for networking (e.g., description documents of digital twin for data centers (Li et al., 2023c)), none of the existing works provide in-depth investigation on whether and how LLMs can be adapted to solve networking tasks. Hence, this work tries to bridge this research gap and proposes an LLM adaptation framework for networking.
@@ 3. Motivation
  • Finally, they involve diverse input modalities, covering the primary data modalities in many networking tasks. For example, many continuous signals in network adaptation problems (e.g., packet loss rate in congestion control (Yen et al., 2023)) are represented as scalar data, while graph representation is widely used in other network system problems like hardware device distributed placement (Mirhoseini et al., 2017).
@@
Challenge 2: Inefficiency of token-based answer generation. As introduced in $\S$2.2, by default, LLM generates answers with LM head by predicting next tokens in an autoregressive manner, which, however, exhibits two main drawbacks in the networking area.
@@
Note that the above problems are not unique to the networking area. Nevertheless, they reduce the efficiency of LLMs in networking, since networking tasks often require high reliability and quick responsiveness (Meng et al., 2020; Chen et al., 2018a).
@@
The adaptation costs become even more expensive when fine-tuning the full parameters of LLMs (Ding et al., 2023; Fu et al., 2023b). As shown in Figure 4, fully fine-tuning Llama2-7B for the VP task consumes 65.88GB GPU memory and 7.9h of training time. This is because full-parameter fine-tune requires extensive memory and computation to store and maintain the training states due to the large parameter size. In fact, another practical drawback associated with full-parameter fine-tune is that it may disrupt the pre-trained knowledge of LLM since it updates all of its parameters. As a result, this may prevent a single LLM to share across different networking tasks.
@@
  • The large modality gap makes the input information of networking tasks incompatible with LLM, preventing the LLM to effectively process task inputs.
  • The default token-based answer generation of LLM exhibits inefficiency in the networking domain, which reduces the reliability and responsiveness for LLM to serve network systems.
  • The large parameter size of LLM leads to significant costs of adapting LLM to acquire domain-specific knowledge for networking, especially for RL-based tasks which require environment interaction.
@@ 4. NetLLM Design
  • Data-driven low-rank networking adaptation (DD-LRNA). To reduce the costs of adaptation, NetLLM develops an efficient DD-LRNA scheme to solve challenge 3. It incorporates a data-driven pipeline to adapt LLM for both prediction and decision-making tasks, and introduces different low-rank matrices to learn domain-specific knowledge to further minimize the adaptation costs.
@@ 4.1. Multimodal Encoder
The key to process task-specific information (e.g., network throughputs) with LLM is to project the multimodal input data into token space to enable efficient utilization by the LLM. To achieve this, we design a multimodal encoder to automatically learn such projection. Figure 6 illustrates the architecture of this module, which incorporates two blocks.
Feature encoder. We first employ different feature encoders to extract features from raw input data of various modalities. A key design consideration here is the choice of feature encoder for each modality. Instead of handcrafting feature encoders from scratch, which entails high model engineering costs, we reuse the existing well-designed encoders tailored for specific modalities. For example, Decima (Mao et al., 2019) for CJS task develops a graph neural network (GNN) (Wu et al., 2021) encoder to extract features from DAG information, and Vision Transformer (ViT) (Dosovitskiy et al., 2021) has been widely used to encode images into hidden features. These designs are precious research outcomes and prove effective in processing specific modalities. Therefore, we cherish and efficiently utilize these designs by integrating them into our multimodal encoder module to encode task input data of diverse modalities.
@@ 4.2. Networking Head
With the multimodal encoder, the LLM is capable to extract high-level features that encompass important task-specific information from input data of various modalities. These features are then fed into the networking head for direct answer generation. Specifically, the networking head is designed as a trainable linear layer to predict task-specific answers based on LLM output features, which can be flexibly customized according to specific networking task. With this head, all answers generated by LLM are guaranteed to be valid, ensuring the reliability of LLM for networking. Moreover, LLM can generate one answer within a single round of inference, thus significantly reducing the generation latency.
Example. Figure 7 compares the difference between LM head for token prediction and networking head with ABR task as the example. As depicted, the LM head generates answers by predicting next tokens autoregressively, which requires multiple rounds of inference and thus entails high generation latency. Besides, due to the inherent uncertainty of token-based prediction, the generated answers may be invalid, such as a bitrate that does not exist. In contrast, the networking head is specially designed to predict the probability distribution of bitrates, enabling direct answer generation within a single round of inference. Furthermore, since the outputs of networking head are limited to the discrete set of candidate bitrates, the generated answers are guaranteed to be always valid. The superiority of networking head over LM head is also illustrated in Figure 2 (middle, right), where NetLLM uses it to ensure the validness of answers and quickly produces answers before the response deadline for VP task.
@@ 4.3. Data-Driven Low-Rank Networking Adaptation
In this part, we delve into the detailed design of the proposed data-driven low-rank networking adaptation (DD-LRNA) scheme to efficiently fine-tune LLM to acquire domain knowledge. The DD-LRNA comprises the following two core designs: i) a data-driven adaptation pipeline for both prediction and decision-making networking tasks; ii) a low-rank adaptation approach to constrain the fine-tune process to a small number of parameters for more efficient adaptation.
Data-Driven networking adaptation. In the case of prediction networking tasks, it is straightforward to fine-tune the LLM through the standard SL data-driven training pipeline. Specifically, given a task-specific dataset $\mathcal{D}_{sl}=\{\mathcal{X},\mathcal{Y}\}$ of inputs $\boldsymbol{x}\in\mathcal{X}$ and labels $y\in\mathcal{Y}$, we leverage the multimodal encoder to encode input data $\boldsymbol{x}$, elicit prediction results $\hat{y}$ from LLM with networking head, and compute loss for parameter update by:
@@
Nevertheless, when it comes to the decision-making tasks, the traditional RL training pipeline becomes impractical due to the time-consuming process of the interaction between LLM and environments. To tackle this challenge, we design our RL adaptation pipeline based on the efficient data-driven RL techniques (Prudencio et al., 2023; Yen et al., 2023), which tackle the same problem as traditional RL but without the need of environment interaction. Specifically, we collect the experience dataset with any existing (non-LLM) policy, and exploit this dataset to fine-tune LLM for reward optimization. Notably, unlike traditional RL that requires periodic refreshing of the experience dataset during training (Van Hasselt et al., 2016; Schulman et al., 2017), our approach allows the dataset to be collected only once and used throughout the entire training process. As a result, the costs of adapting LLM for RL-based networking tasks can be significantly reduced (e.g., with 51.1%/37.7% reduction of training time for ABR/CJS task under the same training iterations, as depicted in Figure 3).
@@
where $w$ is the context window to facilitate the learning of return distribution. Next, we feed data $d$ to the LLM to generate actions $\{\hat{a}_{i}^{1},\cdots,\hat{a}_{i}^{m}\}_{i=t-w+1}^{t}$. Notably, we consider return and each piece of state, action information as different modalities, and process them separately. Finally, the loss for updating parameters is calculated by:
@@
Low-rank networking adaptation. With the data-driven adaptation pipeline in place, the LLM can now be fine-tuned for networking adaptation. Given the pre-trained parameters of LLM denoted as $\Phi_{0}$, the goal of fine-tune is to search for the parameter update $\Delta\Phi$ such that the resulting parameters $\Phi=\Phi_{0}+\Delta\Phi$ are optimized for the specific networking task. Nevertheless, due to the large parameter size of LLM, directly fine-tuning the full parameters of LLM entails prohibitive computation costs, as the dimension of learned parameters $|\Delta\Phi|$ is equal to $|\Phi_{0}|$ (e.g., $|\Phi_{0}|=540$ billion for PaLM (Chowdhery et al., 2023)).
To combat the above limitation, we freeze the parameters of LLM and introduce additional low-rank matrices to approximate the changes needed in the LLM parameters to learn domain-specific knowledge. The underlying insight is that the parameter changes during adaptation (i.e., $\Delta\Phi$) reside on an intrinsic low rank (Aghajanyan et al., 2020; Hu et al., 2021). Therefore, for each pre-trained matrix $W_{0}\in\Phi_{0}$ of dimension $d\times k$, we hypothesize the existence of a low rank $r\ll\min\{d,k\}$ and construct two low-rank matrices $A,B$ of dimension $d\times r,r\times k$ to approximate the update of $W_{0}$, i.e., $W=W_{0}+\Delta W=W_{0}+AB$. During adaptation, $W_{0}$ is frozen and all parameter update is constrained on matrices $A$ and $B$. As shown in Figure 4, this significantly reduces the fine-tune costs with the reduction of 60.9% GPU memory and 15.1% training time, since the low-ranks only introduces 0.31% of trainable parameters. Another benefit of this approach is that the pre-trained knowledge of LLM is preserved as each $W_{0}\in\Phi_{0}$ is retained without any update. Hence, the same LLM can serve as the foundation model for different tasks, and train different copies of $A,B$ to acquire different domain knowledge.
Putting all together. The DD-LRNA scheme is briefly summarized in Figure 8. As shown, we freeze the parameters of LLM and allocate different trainable low-rank matrices for each task. These matrices are then fine-tuned over a dataset to acquire domain-specific knowledge. For decision-making tasks, the dataset is collected by using existing policies to interact with environments. At each fine-tune step, we sample a batch of data from the dataset, feed data to the LLM to generate answers, compute loss according to equation (1) for prediction tasks or equation (4) for decision-making tasks, and update the low-rank matrices through gradient decent.
@@ 4.4. Implementation
NetLLM is fully implemented in Python and Bash, and can be easily integrated into existing SL/RL codebases to adapt LLM for networking tasks. As depicted in Figure 9, it interacts with an existing codebase with three APIs. First, Adapt triggers NetLLM to use the provided dataset to adapt the LLM to learn domain-specific knowledge for the target task, and returns the snapshot of the adapted LLM. Second, Test evaluates the performance of the adapted LLM on the testing environments generated with the given simulation settings. Finally, for RL-based tasks without an available dataset for adaptation, NetLLM offers the RL_Collect to collect the experience dataset by using the given RL policies to interact with the environments. Afterwards, the collected dataset can be plugged to the Adapt API to adapt the LLM.
@@
(a) Average performance with different random seeds
(b) CDF Performance
@@ 5.1. Setup
Metrics. For performance metrics, we consider mean absolute error (MAE) for VP, Quality of Experience (QoE) scores for ABR, job completion time (JCT) for CJS. Lower MAE, higher QoE and lower JCT indicate better performance. In particular, following the same formula in GENET, QoE is calculated as the weighted linear combination of bitrate, rebuffering time and bitrate fluctuation.
@@ 5.2. General Evaluation
[…]
It is worth noting that although the learning-based algorithms yield improvement over traditional rule-based algorithms as depicted in Figure 10, their performance gain relies on engineering specialized DNN models for the target tasks. In contrast, empowered by NetLLM, a single LLM (specifically, Llama2 in this experiment) can serve as the foundation model and share across various tasks, thus significantly reducing the overhead of model engineering.
@@ 5.3. Generalization
From Figure 11, we also notice that learning-based algorithms do not always outperform conventional rule-based algorithms for the ABR task. Figure 12 breaks down the QoE scores of all ABR methods for more detailed analysis. As shown, GENET is surpassed by MPC with 5.2%/5.9% lower average QoE on unseen setting 1/2. More specifically, on unseen setting 1, where the streaming video is different from training one, GENET fails to optimize video bitrates and thus achieves worse performance than MPC. In contrast, NetLLM-adapted Llama2 strikes a good balance between the three QoE factors and thus achieves the highest QoE scores on all settings. These cases exemplify that conventional DNN models may perform poorly in unseen environments. In comparison, by adapting LLM for networking with our NetLLM framework, we can indeed efficiently utilize the extensive knowledge of LLM to achieve stronger generalization.

### v3

@@ 2.3. Domain-Adapted LLMs
The impressive performance of LLMs has sparked pioneering research to adapt LLMs for other domains (Driess et al., 2023; Lin et al., 2023; Liu et al., 2023a; Zitkovich et al., 2023). For example, PaLM-E (Driess et al., 2023) adapts PaLM (Chowdhery et al., 2023) to generate step-by-step controlling commands for robotic manipulation. ESMFFold (Lin et al., 2023) showcases the successful applications of LLM in biological fields, which leverages LLM to predict atomic-level protein structure.
Inspired by these promising outcomes, this work explores the adaptation of LLMs for networking to address the limitations of existing learning-based algorithms, hoping to pave the way for more sustainable design philosophy of networking solutions. Unfortunately, although some prior studies have showcased that LLMs are capable of generating some technical documents for networking (e.g., description documents of digital twin for data centers (Li et al., 2023d)), none of the existing works provide an in-depth investigation on whether and how LLMs can be adapted to solve networking tasks. Hence, this work tries to bridge this research gap and proposes an LLM adaptation framework for networking.
Figure 1. Illustration of the token-based answer generation mechanism of LLMs.
@@ 3. Motivation
  • Finally, they involve diverse input modalities, covering the primary data modalities in many networking tasks. For example, many continuous signals in network adaptation problems (e.g., packet loss rate in congestion control (Yen et al., 2023)) are represented as scalar data.
@@
Challenge 2: Inefficiency of token-based answer generation. As introduced in $\S$2.2, by default, LLMs generate answers with LM head by predicting next tokens in an autoregressive manner. This, however, exhibits two main drawbacks in the networking area.
@@
Note that the above problems are not unique to the networking area. Nevertheless, they reduce the efficiency of LLMs in networking, since networking tasks often require high reliability and quick responsiveness (Meng et al., 2020; Chen et al., 2018).
@@
The adaptation costs become even more expensive when fine-tuning the full parameters of LLMs (Ding et al., 2023; Fu et al., 2023b). As shown in Figure 4, fully fine-tuning Llama2-7B for the VP task consumes 65.88GB GPU memory and 7.9h of training time. This is because full-parameter fine-tune requires extensive memory and computation to store and maintain the training states due to the large parameter size. In fact, another practical drawback associated with full-parameter fine-tune is that it may disrupt the pre-trained knowledge of LLM since it updates all the parameters of LLM. As a result, this may prevent a single LLM to share across different networking tasks.
@@
  • The large modality gap makes the input information of networking tasks incompatible with the LLM, preventing the LLM from effectively processing task inputs.
  • The default token-based answer generation of the LLM exhibits inefficiency in the networking domain, which reduces the reliability and responsiveness for the LLM to serve network systems.
  • The large parameter size of the LLM leads to significant costs to acquire domain-specific knowledge for networking, especially for RL-based tasks which require environment interaction.
@@ 4. NetLLM Design
  • Data-driven low-rank networking adaptation (DD-LRNA). To reduce the costs of adaptation, NetLLM develops an efficient DD-LRNA scheme to solve challenge 3. It incorporates a data-driven pipeline to adapt LLMs for both prediction and decision-making tasks, and introduces different low-rank matrices to learn domain knowledge to further minimize the adaptation costs.
@@ 4.1. Multimodal Encoder
The key to processing task-specific information is to project the multimodal input data into token space to enable efficient utilization by the LLM. To achieve this, we design a multimodal encoder to automatically learn such projection. Figure 6 illustrates the architecture of this module, which incorporates two blocks.
Feature encoder. We first employ different feature encoders to extract features from raw input data of various modalities. A key design consideration here is the choice of feature encoder for each modality. Instead of handcrafting feature encoders from scratch, which entails high model engineering costs, we reuse the existing well-designed encoders tailored for specific modalities. For example, Decima (Mao et al., 2019) for CJS task develops a graph neural network (GNN) (Wu et al., 2021) encoder to extract features from DAG information, and Vision Transformer (ViT) (Dosovitskiy et al., 2021) has been widely used to encode images into hidden features. These designs are precious research outcomes and prove effective in processing specific modalities. Therefore, we cherish and efficiently utilize these designs by integrating them into our multimodal encoder module. Specifically, we leverage the following feature encoders to encode different data modalities involved in networking: ViT for images, 1D-CNN for time-series and sequence data (e.g., historical throughputs and future chunk sizes at different bitrates in ABR), fully connected layer for scalar data (e.g., buffer occupancy in ABR), and GNN for graph information (e.g., DAGs in CJS).
@@ 4.2. Networking Head
With the multimodal encoder, the LLM is capable to extract high-level features that encompass important task-specific information from input data of various modalities. These features are then fed into the networking head for direct answer generation. Specifically, the networking head is designed as a trainable linear layer to predict task-specific answers based on LLM output features, which can be flexibly customized according to specific networking tasks. Unlike LM head, the networking head constrains the answer generation into the valid range of possible answers, such as the range of valid viewport coordinates in VP and the set of video bitrates in ABR. In this way, all answers generated by LLM are guaranteed to be valid, ensuring the reliability of LLM for networking. Moreover, LLM can generate one answer within a single round of inference, thus significantly reducing the generation latency.
Example. Figure 7 compares the difference between LM head for token prediction and networking head with ABR task as the example. As depicted, the LM head generates answers by predicting next tokens autoregressively, which requires multiple rounds of inference and thus entails high generation latency. Besides, due to the inherent uncertainty of token-based prediction, the generated answers may be invalid, such as a bitrate that does not exist. In contrast, the networking head is specially designed to predict the probability distribution of bitrates, enabling direct answer generation within a single round of inference (e.g., the bitrate with the maximum probability is chosen as the answer in Figure 7). Furthermore, since the outputs of networking head are limited to the discrete set of candidate bitrates, the generated answers are guaranteed to be always valid. The superiority of networking head over LM head is also illustrated in Figure 2 (middle, right), where NetLLM uses it to ensure the validness of answers and quickly produces answers before the response deadline for VP task.
@@ 4.3. Data-Driven Low-Rank Networking Adaptation
In this part, we delve into the detailed design of the proposed data-driven low-rank networking adaptation (DD-LRNA) scheme to efficiently fine-tune LLMs to acquire domain knowledge. The DD-LRNA comprises the following two core designs: i) a data-driven adaptation pipeline for both prediction and decision-making networking tasks; ii) a low-rank adaptation approach to constrain the fine-tuning process to a small number of parameters for more efficient adaptation.
Data-driven networking adaptation. In the case of prediction networking tasks, it is straightforward to fine-tune the LLM through the standard SL data-driven training pipeline. Specifically, given a task-specific dataset $\mathcal{D}_{sl}=\{\mathcal{X},\mathcal{Y}\}$ of inputs $\boldsymbol{x}\in\mathcal{X}$ and labels $y\in\mathcal{Y}$, we leverage the multimodal encoder to encode input data $\boldsymbol{x}$, elicit prediction results $\hat{y}$ from the LLM with networking head, and compute loss for parameter update by:
@@
Nevertheless, when it comes to the decision-making tasks, the traditional RL training pipeline becomes impractical due to the time-consuming process of the interaction between the LLM and environments. To tackle this challenge, we design our RL adaptation pipeline based on the efficient data-driven RL techniques (Prudencio et al., 2023; Yen et al., 2023), which tackle the same problem as traditional RL but without the need of environment interaction. Specifically, we collect the experience dataset with any existing (non-LLM) networking algorithms, and exploit this dataset to fine-tune the LLM for reward optimization. The intuition behind is to let the LLM learn a better policy by observing the behaviors of existing algorithms (Levine et al., 2020). That means, the LLM should not only learn from the good actions that lead to good performance, but also try to understand why some actions will lead to bad performance (Yen et al., 2023).
It is worth noting that, unlike traditional RL that requires periodic refreshing of the experience dataset during training (Van Hasselt et al., 2016; Schulman et al., 2017), our approach allows the dataset to be collected only once and used throughout the entire training process. As a result, the costs of adapting LLMs for RL-based networking tasks can be significantly reduced (e.g., with 51.1%/37.7% reduction of training time for ABR/CJS task under the same training iterations, as depicted in Figure 3).
@@
where $w$ is the context window to facilitate the learning of return distribution. Next, we feed data $d$ to the LLM to generate actions $\{\hat{a}_{i}^{1},\cdots,\hat{a}_{i}^{m}\}_{i=t-w+1}^{t}$. Notably, we consider return and each piece of state, action information as different modalities, and process them separately. Finally, the training loss is calculated by:
@@
Low-rank networking adaptation. With the data-driven adaptation pipeline in place, the LLM can now be fine-tuned for networking adaptation. Given the pre-trained parameters of the LLM denoted as $\Phi_{0}$, the goal of fine-tuning is to search for the parameter update $\Delta\Phi$ such that the resulting parameters $\Phi=\Phi_{0}+\Delta\Phi$ are optimized for the specific networking task. Nevertheless, due to the large parameter size of the LLM, directly fine-tuning the full parameters entails prohibitive computation costs, as the dimension of learned parameters $|\Delta\Phi|$ is equal to $|\Phi_{0}|$ (e.g., $|\Phi_{0}|=540$ billion for PaLM (Chowdhery et al., 2023)).
To combat the above limitation, we freeze the parameters of LLM and introduce additional low-rank matrices to approximate the changes needed in the LLM parameters to learn domain-specific knowledge. The underlying insight is that the parameter changes during adaptation (i.e., $\Delta\Phi$) reside on an intrinsic low rank (Aghajanyan et al., 2020; Hu et al., 2021). Therefore, for each pre-trained matrix $W_{0}\in\Phi_{0}$ of dimension $d\times k$, we hypothesize the existence of a low rank $r\ll\min\{d,k\}$ and construct two low-rank matrices $A,B$ of dimension $d\times r,r\times k$ to approximate the update of $W_{0}$, i.e., $W=W_{0}+\Delta W=W_{0}+AB$. During adaptation, $W_{0}$ is frozen and all parameter update is constrained on matrices $A$ and $B$. As shown in Figure 4, this significantly reduces the fine-tuning costs with the reduction of 60.9% GPU memory and 15.1% training time, since the low-ranks only introduces 0.31% of trainable parameters. Another benefit of this approach is that the pre-trained knowledge of the LLM is preserved as each $W_{0}\in\Phi_{0}$ is retained without any update. Hence, the same LLM can serve as the foundation model for different tasks, and train different copies of $A,B$ to acquire different domain knowledge.
Putting all together. The DD-LRNA scheme is briefly summarized in Figure 8. As shown, we freeze the parameters of LLM and allocate different trainable low-rank matrices for each task. These matrices are then fine-tuned over a dataset to acquire domain-specific knowledge. For decision-making tasks, the dataset is collected by using existing algorithms to interact with environments. At each fine-tuning step, we sample a batch of data from the dataset, feed data to the LLM to generate answers, compute loss according to equation (1) for prediction tasks or equation (4) for decision-making tasks, and update the low-rank matrices through gradient descent. Note that in addition to low-rank matrices, the gradients are also propagated to update the parameters of multimodal encoder and networking head for performance optimization.
@@ 4.4. Implementation
NetLLM is fully implemented in Python and Bash, and can be easily integrated into existing SL/RL codebases to adapt LLMs for networking tasks. As depicted in Figure 9, it interacts with an existing codebase with three APIs. First, Adapt triggers NetLLM to use the provided dataset to adapt the LLM to learn domain-specific knowledge for the target task, and returns the snapshot of the adapted LLM. Second, Test evaluates the performance of the adapted LLM on the testing environments generated with the given simulation settings. Finally, for RL-based tasks without an available dataset for adaptation, NetLLM offers the RL_Collect API to collect the experience dataset by using the given RL policies to interact with the environments. Afterwards, the collected dataset can be plugged into the Adapt API to adapt the LLM.
@@ 5.1. Setup
[…]
@@
(a) Average performance with different random seeds
(b) CDF Performance
5.2. General Evaluation
[…]
@@ 5.2. General Evaluation
It can be seen that there is a clear ranking among the three baselines on each task, where the learning-based algorithms consistently yield improvement over traditional rule-based algorithms. This can be attributed to the inherent strength of DNNs in fitting complex functions for prediction. However, the LLM demonstrates even more powerful capabilities in function approximation, pattern mining and long-term planning, owing to its large parameter size and large-scale pre-training. Hence, by effectively utilizing the strengths of LLM to solve networking tasks, NetLLM achieves superior performance compared to the other learning-based algorithms. In addition, it is worth mentioning that the performance gain of learning-based algorithms relies on engineering specialized DNN models for the target tasks. In contrast, NetLLM efficiently utilizes the LLM as the foundation model for task solving. That said, it uses the same LLM to solve various networking tasks without any further modification on the model, thus significantly reducing the overhead of model engineering.
@@ 5.3. Generalization
From Figure 11, we also notice that learning-based algorithms do not always outperform conventional rule-based algorithms for the ABR task. Figure 12 breaks down the QoE scores of all ABR methods for more detailed analysis. As shown, GENET is surpassed by MPC with 5.2%/5.9% lower average QoE on unseen setting 1/2. More specifically, on unseen setting 1, where the streaming video is different from training one, GENET fails to optimize video bitrates and thus achieves worse performance than MPC. On the other hand, GENET struggles to adapt to the dynamic fluctuations of bandwidth on unseen setting 2, where the testing bandwidth traces change more frequently than training ones. It may inappropriately select high bitrates when the current bandwidth resources become scarce, and thus produces the highest rebuffering time among other methods. In contrast, NetLLM-adapted Llama2 strikes a good balance between the three QoE factors and thus achieves the highest QoE scores on all settings. These cases exemplify that conventional DNN models may perform poorly in unseen environments. In comparison, leveraging our NetLLM framework, we can indeed efficiently utilize the extensive knowledge of the LLM to achieve stronger generalization.